# ResSVD: Residual Compensated SVD for Large Language Model Compression

### v2

@@ 3.1 Residual Compensation for SVD Truncation
The direct truncation loss $\mathcal{L}_{D}$ between $\boldsymbol{W}$ and $\boldsymbol{W}_{r}$ is given by:
$\mathcal{L}_{D}=\big{\|}\boldsymbol{W}-\boldsymbol{W}_{r}\big{\|}_{F}^{2}=\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}(\boldsymbol{W}).$ (9)
According to Lemma 3.2, the compensated truncation loss $\mathcal{L}_{C}$, as defined in Equation (7), is upper-bounded by:
$\sum_{i=r_{2}+1}^{rank(\boldsymbol{R})}\sigma_{i}^{2}(\boldsymbol{R})\leq\sum_{i=r_{1}+r_{2}+1}^{\min(m,n)}\sigma_{i}^{2}(\boldsymbol{W})=\sum_{i=r+1}^{\min(m,n)}\sigma_{i}^{2}(\boldsymbol{W}).$ (10)
Therefore, we obtain:
$\boxed{\mathcal{L}_{C}\leq\mathcal{L}_{D}}$ (11)
This proves that the compensated truncation loss $\mathcal{L}_{C}$ is smaller than the direct truncation loss $\mathcal{L}_{D}$, which means that $\hat{\boldsymbol{W}}_{r}$ is closer to the original weight matrix $\boldsymbol{W}$. ∎
@@ 3.2 Partial-layer Compression for SVD
Previous works [15, 44, 41, 22] compress all model layers even if they assign layer-specific ratios based on their relative importance, which often leads to a high layer-wise error, resulting in noticeable degradation in the performance of compressed models. We compare the layer-wise error of LLMs across four families with different layer selection strategies, the results are shown in Figure 2. There is a significant error propagation across multiple LLM families, where the error progressively accumulates layer by layer during the forward pass. This phenomenon is particularly pronounced when compressing the first 8 layers. To satisfy the overall compression ratio, a higher layer compression ratio (80% in this case) must be applied to these layers. Compressing the model at such an early stage introduces substantial approximation errors, which are then propagated through the remaining layers. Although the later layers are left uncompressed, the forward pass carries the accumulated error through the network. Consequently, in all evaluated LLM families, these early compressed layers exhibit the highest layer-wise error across model layers. In contrast, when only compressing the last few layers, the earlier layers remain untouched with zero error. Although the compression ratio for selected layers is relatively high, leading to slightly faster error accumulation, the overall error remains significantly lower than that of compressing all layers.
Interestingly, we observe that although the final-layer error converges to a narrow range regardless of how many of the last layers are compressed, the differences within this range still exhibit a strong influence on model performance. As illustrated in the top figure of Figure 3, there is a strong correlation between the final-layer error and the average zero-shot accuracy. Motivated by this observation, we select the number of last layers to compress by minimizing the final-layer error, and the layer compression ratio $R_{o}$ satisfies:
@@
where $N$ is the number of model layers, $R_{l}$ represents the overall compression ratio. As shown in the bottom figure of Figure 3, compared to SVD-LLM [41], our strategy significantly reduces layer-wise error across all model layers and mitigates error propagation.
Figure 2: Layer-wise error comparison between original model and multiple LLM structures compressed by ResSVD with different layer selection strategies on WikiText-2 [29]. The overall compression ratio is 20%, and different layer selection strategies strictly adhere to the overall compression ratio constraint.
Figure 3: (Top) Kendall correlation. (Bottom) Layer-wise error comparison between ResSVD and SVD-LLM [41] under 20% compression ratio.
@@ 4.1 Setups
Baselines. We compare with four baselines in post-training setting, i.e., without re-training of the model parameters: conventional SVD, ASVD [44], SVD-LLM [41], and AdaSVD [22].
Models and Datasets. We evaluate ResSVD on ten models spanning four LLM families: LLaMA-7B [38], LLaMA-13B [38], LLaMA-30B [38], LLaMA-2-7B [39], LLaMA-2-13B [39], OPT-6.7B [46], OPT-13B [46], OPT-30B [46], Mistral-7B [17], Vicuna-7B [5]. For language modeling, we use three benchmark datasets: WikiText-2 [29], PTB [28], and C4 [32]. For zero-shot reasoning and understanding, we evaluate on seven tasks within the LM-Evaluation-Harness framework [^1] https://github.com/EleutherAI/lm-evaluation-harness: OpenbookQA [30], WinoGrande [34], HellaSwag [45], PIQA [3], MathQA [1], ARC_e and ARC_c [7].
Implementation Details.   To facilitate a fair comparison, we follow the protocols of ASVD, SVD-LLM, and AdaSVD, randomly selecting 256 calibration samples with a sequence length of 2048 from WikiText-2. We then apply data whitening prior to performing SVD truncation. All results are reproduced by re-running their respective open-source codebases, except for AdaSVD, whose results are directly taken from the original paper due to the lack of released code. All methods are all implemented with PyTorch [^2] https://github.com/pytorch/pytorchandHuggingFace and Transformers [^3] https://github.com/huggingface/transformers on NVIDIA A100 GPUs.
@@ 4.2 Results
We evaluate the overall performance of ResSVD from three aspects: ① Effectiveness under different compression ratios (ranging from 20% to 60% in increments of 10%), ② Scalability when applied to larger-scale models, and ③ Generalizability across diverse LLM families. In addition, qualitative results, such as generated contents produced by compressed models, are provided in Appendix A.4, offering a more intuitive and direct comparison.
Table 1: Overall performance of LLaMA-7B [38] compressed by ResSVD and baselines under 20% to 60% compression ratios ("Ratio"), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\uparrow$)). The best results are marked in bold. Blue arrows within parentheses highlight the relative improvement over the second-best method.
@@
- Original 5.67 8.80 7.63 0.28 0.67 0.67 0.56 0.38 0.78 0.27 0.52
20% SVD 20082.86 20338.96 18784.20 0.14 0.27 0.51 0.26 0.21 0.53 0.21 0.30
ASVD [44] 9.27 15.09 13.68 0.25 0.53 0.60 0.41 0.27 0.68 0.25 0.43
SVD-LLM [41] 7.89 16.54 15.92 0.23 0.56 0.62 0.42 0.29 0.69 0.23 0.43
ResSVD 7.47 ($\downarrow$5%) 12.27 ($\downarrow$19%) 12.22 ($\downarrow$11%) 0.25 0.59 0.67 0.47 0.34 0.71 0.25 0.47 ($\uparrow$9%)
30% SVD 13155.97 17354.46 21012.91 0.13 0.26 0.51 0.25 0.21 0.54 0.21 0.30
ASVD [44] 222.98 586.79 148.79 0.15 0.32 0.53 0.30 0.21 0.59 0.21 0.33
SVD-LLM [41] 9.52 28.97 26.38 0.20 0.49 0.59 0.37 0.27 0.65 0.22 0.40
ResSVD 9.52 (-) 20.31 ($\downarrow$30%) 18.29 ($\downarrow$31%) 0.23 0.54 0.63 0.41 0.30 0.67 0.23 0.43 ($\uparrow$8%)
40% SVD 52326.99 59859.41 47643.04 0.15 0.25 0.51 0.25 0.21 0.52 0.20 0.30
ASVD [44] 5262.11 8806.33 6522.61 0.14 0.26 0.49 0.26 0.22 0.53 0.20 0.30
SVD-LLM [41] 13.83 57.07 48.47 0.19 0.41 0.58 0.32 0.22 0.59 0.22 0.36
ResSVD 12.92 ($\downarrow$7%) 46.93 ($\downarrow$18%) 30.51 ($\downarrow$37%) 0.19 0.44 0.60 0.35 0.27 0.62 0.24 0.39 ($\uparrow$8%)
50% SVD 130388.72 86721.38 79853.46 0.16 0.26 0.49 0.25 0.22 0.52 0.18 0.30
ASVD [44] 62726.86 117959.06 77773.84 0.13 0.25 0.48 0.25 0.22 0.53 0.21 0.30
SVD-LLM [41] 24.05 150.58 141.87 0.16 0.34 0.55 0.29 0.21 0.56 0.22 0.33
ResSVD 22.41 ($\downarrow$7%) 128.80 ($\downarrow$14%) 83.80 ($\downarrow$41%) 0.16 0.36 0.58 0.32 0.23 0.59 0.22 0.35 ($\uparrow$6%)
60% SVD 52326.99 59859.41 47643.04 0.15 0.25 0.51 0.25 0.21 0.52 0.20 0.30
ASVD [44] 16221.43 20119.36 16561.39 0.13 0.26 0.50 0.25 0.23 0.53 0.18 0.30
SVD-LLM [41] 53.20 378.19 310.17 0.12 0.29 0.52 0.28 0.20 0.55 0.22 0.31
ResSVD 48.67 ($\downarrow$9%) 323.74 ($\downarrow$14%) 260.06 ($\downarrow$16%) 0.14 0.30 0.53 0.28 0.21 0.55 0.21 0.32 ($\uparrow$3%)
@@
Performance under different compression ratios. We evaluate the performance of LLaMA-7B [38] and LLaMA-2-7B [39] compressed by ResSVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across 10 benchmark datasets. The results for LLaMA-7B are presented in Table 1, while those for LLaMA-2-7B are provided in Appendix A.2. On the three language modeling datasets, WikiText-2 [29], PTB [28], and C4 [32], ResSVD consistently outperforms all baselines across evaluated compression ratios. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ResSVD exhibits stronger generalization capability. More importantly, under a relatively high compression ratio (e.g., 50%), ResSVD still achieves substantial perplexity reductions compared to the existing best-performing baseline SVD-LLM [41]: 7% on WikiText-2, 14% on PTB, and 41% on C4. This demonstrates that ResSVD maintains an obvious performance advantage even under aggressive compression. In addition, on seven common sense reasoning datasets, ResSVD surpasses the existing best baseline on the majority of datasets, with an average accuracy gain of up to 9%, and a minimum improvement of 3% improvement, further highlighting its robustness and overall effectiveness.
Performance on Larger Scale LLMs. To examine the scalability and robustness of ResSVD, we evaluate its performance on larger scale LLMs from two representative LLM families: LLaMA (including LLaMA-13B, LLaMA-30B, and LLaMA-2-13B) and OPT (OPT-13B and OPT-30B). As presented in Table 2, ResSVD consistently achieves superior performance over existing baselines under 20% compression ratio, demonstrating robust effectiveness across varying model scales.
@@
Performance on Different LLM families. To evaluate the generalization ability of ResSVD, we apply it to four LLMs from different families, including OPT-6.7B [46], LLaMA-2-7B, Mistral-7B [17], and Vicuna-7B [5]. As shown in Table 3, under 30% compression ratio, ResSVD consistently outperforms all baselines across three language modeling benchmarks across these diverse architectures. The most significant relative improvement reaches 75% on LLaMA-2-7B. Moreover, the best overall improvement is achieved on Mistral-7B, with perplexity reductions of 71% on WikiText-2, 45% on PTB, and 46% on C4. We reproduce ASVD [44] and SVD-LLM [41] using their public codebases. While ASVD fails in certain cases due to numerical instability (denoted as NaN in the table), ResSVD consistently maintains stable and reliable performance.
Table 2: Perplexity ($\downarrow$) of larger-scale LLMs under 20% compression ratio on PTB [28].
Method LLaMA-13B LLaMA-30B LLaMA-2-13B OPT-13B OPT-30B
SVD 1878.04 555.55 5464.57 1552.55 250.49
ASVD [44] 12.42 31.36 81.00 36.47 30.95
SVD-LLM [41] 12.17 9.10 88.13 14.86 12.94
ResSVD 9.70 8.41 66.47 13.22 12.89
@@
ASVD [44] 26.67 71.36 44.51
SVD-LLM [41] 28.03 37.46 40.35
ResSVD 17.10 ($\downarrow$36%) 27.24 ($\downarrow$27%) 38.40 ($\downarrow$5%)
LLaMA-2-7B SVD 30373.39 48930.94 36905.54
ASVD [44] 984.03 NaN NaN
SVD-LLM [41] 10.66 292.90 34.96
ResSVD 10.32 ($\downarrow$3%) 73.04 ($\downarrow$75%) 23.28 ($\downarrow$33%)
@@
ASVD [44] 221.66 927.15 266.04
SVD-LLM [41] 48.94 193.22 56.55
ResSVD 14.09 ($\downarrow$71%) 105.37 ($\downarrow$45%) 30.72 ($\downarrow$46%)
@@
ASVD [44] 106.32 NaN NaN
SVD-LLM [41] 12.42 104.27 39.55
ResSVD 11.57 ($\downarrow$7%) 69.28 ($\downarrow$34%) 27.24 ($\downarrow$31%)

### v3

@@ 3.1 Residual Compensation for SVD Truncation
The direct truncation method extracts the first $r$ components from the $\boldsymbol{WS}$ at once. We can decompose it into two parts: the first $r_{i}$ components and the last $r_{r}$ components:
$W_{r}=\underbrace{SVD_{r_{i}}(\boldsymbol{WS})\boldsymbol{S}^{-1}}_{\boldsymbol{W}_{r_{i}}}+\underbrace{(SVD_{r}(\boldsymbol{WS})-SVD_{r_{i}}(\boldsymbol{WS}))\boldsymbol{S}^{-1}}_{\Delta}$ (8)
For ERC-SVD, the truncation loss is given by:
$\big\|\boldsymbol{W}-\hat{\boldsymbol{W}}_{r}\big\|_{F}^{2}=\big\|\boldsymbol{W}-(\boldsymbol{W}_{r_{i}}+\boldsymbol{R}_{r_{r}})\big\|_{F}^{2}=\big\|\boldsymbol{R}-\boldsymbol{R}_{r_{r}}\big\|_{F}^{2}.$ (9)
Given that $\boldsymbol{R}_{r_{r}}$ is obtained by direct SVD, which is the optimal rank-$r_{r}$ approximation of $\boldsymbol{R}$. According to Lemma 1, for any matrix $\boldsymbol{X}$ with rank $r_{r}$, we have:
$\big\|\boldsymbol{R}-\boldsymbol{R}_{r_{r}}\big\|_{F}^{2}\leq\big\|\boldsymbol{R}-\boldsymbol{X}\big\|_{F}^{2}.$ (10)
For the direct truncation method, the truncation loss is given by:
$\big\|\boldsymbol{W}-\boldsymbol{W}_{r}\big\|_{F}^{2}=\big\|\boldsymbol{W}-(\boldsymbol{W}_{r_{i}}+\Delta)\big\|_{F}^{2}=\big\|\boldsymbol{R}-\Delta\big\|_{F}^{2}.$ (11)
$\boldsymbol{R}_{r_{r}}$ is the optimal rank-$r_{r}$ approximation, whereas $\Delta$ is an approximation derived from the weighted space $\boldsymbol{WS}$. Hence, by combining Lemma 1 and Inequality 10, the following relationship holds:
$\big\|\boldsymbol{R}-\boldsymbol{R}_{r_{r}}\big\|_{F}^{2}\leq\big\|\boldsymbol{R}-\Delta\big\|_{F}^{2}.$ (12)
By substituting these results into the original form, we obtain the final inequality:
$\boxed{\big\|\boldsymbol{W}-\hat{\boldsymbol{W}}_{r}\big\|_{F}^{2}\leq\big\|\boldsymbol{W}-\boldsymbol{W}_{r}\big\|_{F}^{2}}$ (13)
Therefore, the residual compensation strategy provides a closer approximation to the original weight matrix $\boldsymbol{W}$ than direct truncation methods. ∎
@@ 3.2 Partial-layer Compression for SVD
Previous works [yuan2023asvd, wang2024svd, wang2024basis, li2025adasvd] compress all model layers even if they assign layer-specific ratios based on their relative importance, which often leads to a high layer-wise error, resulting in noticeable performance degradation. We compare the layer-wise error of LLMs across four families with different layer selection strategies; the results of LLaMA-7B and OPT-6.7B are shown in Figure 3, the results of other models are presented in Figure 4. There is a significant error propagation across multiple LLM families, where the error progressively accumulates layer by layer during the forward pass. This phenomenon is particularly pronounced when compressing only the first few layers. A higher layer compression ratio must be applied to these layers, with the constraint of the overall compression ratio. Compressing the model at such an early stage introduces substantial approximation errors, which are then propagated through the remaining layers. Although the later layers are left uncompressed, the forward pass carries the accumulated error through the network. Consequently, in all evaluated LLM families, these early compressed layers exhibit the highest layer-wise error across model layers. In contrast, when only compressing the last few layers, the earlier layers introduce zero error. Although the relatively high compression ratio leads to slightly faster error accumulation, the overall error remains significantly lower than that of compressing all layers.
Interestingly, although the final-layer error converges to a narrow range regardless of how many the last layers are compressed, differences within this range still affect model performance. As illustrated in the left figure of Figure 1, there is a strong correlation between the final-layer error and the average zero-shot accuracy. Motivated by this observation, we select the number of last layers to compress by minimizing the final-layer error, and the layer compression ratio $R_{l}$ must satisfy the following mathematical relation to ensure a fair comparison:
@@
where $N$ is the number of model layers, $R_{o}$ represents the overall compression ratio. As shown in the middle figure of Figure 1, compared to SVD-LLM [wang2024svd], our strategy significantly reduces layer-wise error across all model layers and mitigates error propagation.
Figure 3: Layer-wise error comparison between the original model, LLaMA-7B, and OPT-6.7B compressed by ERC-SVD with different layer selection strategies on WikiText-2. The overall compression ratio is 20%, and all layer selection strategies strictly adhere to the compression constraint.
@@ 4.1 Setups
Evaluation Details.
We compare our method against five baselines without re-training: conventional SVD, ASVD [yuan2023asvd], SVD-LLM [wang2024svd], Basis Sharing [wang2024basis], and AdaSVD [li2025adasvd]. We evaluate ERC-SVD on ten models spanning five LLM families: LLaMA-2-7B [touvron2023llama2], LLaMA-13B [touvron2023llama], LLaMA-2-13B [touvron2023llama2], LLaMA-30B [touvron2023llama], OPT-6.7B [zhang2022opt], OPT-13B [zhang2022opt], OPT-30B [zhang2022opt], Mistral-7B [jiang2023mistral], Vicuna-7B [chiang2023vicuna], Qwen-3-8B [yang2025qwen3]. For language modeling, we use three benchmark datasets: WikiText-2 [merity2016pointer], PTB [marcus1993building], and C4 [raffel2020exploring]. For zero-shot reasoning and understanding, we evaluate on seven tasks within the LM-Evaluation-Harness framework: OpenbookQA [mihaylov2018can], WinoGrande [sakaguchi2021winogrande], HellaSwag [zellers2019hellaswag], PIQA [bisk2020piqa], MathQA [amini2019mathqa], ARC_e and ARC_c [clark2018think]. Moreover, to demonstrate the generality and effectiveness of ERC-SVD, we further extend our evaluation to vision-language models (VLMs), LLaVA [liu2024improved].
Implementation Details.
We follow the protocols of ASVD, SVD-LLM, and Basis Sharing, randomly selecting 256 calibration samples with a sequence length of 2048 from WikiText-2. We then apply matrix scaling prior to performing SVD truncation. All methods are implemented with PyTorch and Transformers on NVIDIA A100 GPUs.
@@ 4.2 Results
We evaluate the overall performance of ERC-SVD across four dimensions: ① Effectiveness under different compression ratios (ranging from 20% to 60%), ② Scalability to larger-scale models, ③ Generalizability across diverse LLM families, and ④ Performance on VLMs, including quantitative and qualitative results, are provided in Table 14 and Appendix A.8, respectively.
Table 1: Overall performance of LLaMA-2-7B [touvron2023llama2] compressed by ERC-SVD and baselines under 20% to 60% compression ratios (“Ratio”), including performance on three language modeling datasets (measured by perplexity ($\downarrow$)) and zero-shot performance on seven common sense reasoning datasets (measured by individual and average accuracy ($\uparrow$)). The best results are marked in bold. NaN denotes evaluation failure due to numerical instability. * refers to results derived from the original paper. - means that results are not available.
@@
- Original 5.47 26.84 7.28 0.31 0.69 0.67 0.56 0.40 0.78 0.27 0.53
ASVD 9.56 120.74 12.85 0.25 0.59 0.61 0.46 0.32 0.72 0.24 0.45
SVD-LLM 8.37 139.68 20.18 0.23 0.50 0.59 0.39 0.26 0.65 0.23 0.41
Basis Sharing 7.77 60.00 16.22 0.24 0.60 0.61 0.43 0.31 0.69 0.24 0.45
ERC-SVD 7.63 45.37 14.73 0.28 0.61 0.65 0.50 0.35 0.73 0.26 0.48
ASVD 984.03 NaN NaN 0.15 0.27 0.51 0.26 0.22 0.53 0.20 0.31
SVD-LLM 10.66 292.90 34.96 0.21 0.42 0.55 0.34 0.22 0.60 0.23 0.37
Basis Sharing 9.69 101.22 25.45 0.23 0.54 0.59 0.38 0.27 0.64 0.23 0.41
ERC-SVD 10.32 73.04 23.28 0.23 0.51 0.62 0.42 0.29 0.68 0.24 0.43
ASVD NaN NaN NaN 0.15 0.25 0.50 0.26 0.22 0.52 0.18 0.30
SVD-LLM 16.11 717.34 61.96 0.16 0.35 0.55 0.30 0.20 0.57 0.23 0.34
Basis Sharing 14.66 207.49 47.41 0.18 0.43 0.57 0.33 0.22 0.61 0.23 0.37
AdaSVD 14.76 304.62 56.98 0.19 0.41 0.58 0.32 0.23 0.58 - -
ERC-SVD 14.17 136.32 43.19 0.20 0.43 0.57 0.35 0.24 0.63 0.23 0.38
ASVD NaN NaN NaN 0.13 0.26 0.50 0.25 0.23 0.50 0.20 0.30
SVD-LLM 27.19 1775.52 129.71 0.14 0.30 0.50 0.28 0.20 0.54 0.23 0.31
Basis Sharing 23.78 689.03 115.84 0.15 0.34 0.54 0.29 0.20 0.56 0.23 0.33
AdaSVD 25.58 593.14 113.84 0.15 0.34 0.54 0.29 0.20 0.56 - -
ERC-SVD 24.26 286.24 100.34 0.14 0.35 0.55 0.31 0.22 0.59 0.22 0.34
ASVD NaN 19581.17 NaN 0.15 0.25 0.50 0.25 0.23 0.52 0.12 0.29
SVD-LLM 54.19 3442.74 263.02 0.14 0.26 0.50 0.27 0.20 0.53 0.21 0.30
Basis Sharing 50.59 2055.63 299.36 0.14 0.29 0.51 0.27 0.20 0.52 0.21 0.30
AdaSVD 60.08 2137.28 294.26 0.12 0.27 0.50 0.27 0.20 0.53 - -
ERC-SVD 68.59 991.48 255.70 0.13 0.29 0.52 0.28 0.21 0.55 0.22 0.31
Performance under different compression ratios.
@@
We evaluate the performance of LLaMA-2-7B [touvron2023llama2] compressed by ERC-SVD, conventional SVD, and existing post-training baselines under compression ratios ranging from 20% to 60% across ten benchmark datasets. The results for LLaMA-2-7B are presented in Table 1. Across three language modeling datasets, WikiText-2 [merity2016pointer], PTB [marcus1993building], and C4 [raffel2020exploring], ERC-SVD consistently outperforms all baselines across most evaluated compression ratios, with only slight suboptimalities observed in specific cases. In particular, on PTB and C4, the improvements are more pronounced, suggesting that ERC-SVD exhibits stronger generalization capability. In addition, across seven common sense reasoning datasets, ERC-SVD outperforms the strongest existing baseline on most datasets, achieving the highest average accuracy across all evaluated compression ratios and further demonstrating its robustness and overall effectiveness.
Performance on Multiple LLM families.
@@
To evaluate the generalization ability of ERC-SVD, we apply it to three LLMs from different families, including OPT-6.7B [zhang2022opt], Mistral-7B [jiang2023mistral], and Vicuna-7B [chiang2023vicuna]. As shown in Table 2, under 30% compression ratio, ERC-SVD consistently outperforms all baselines on three language modeling benchmarks across these architectures. Notably, the largest overall improvements are observed on Mistral-7B, with perplexity reductions of 71% on WikiText-2, 45% on PTB, and 46% on C4. We reproduce ASVD and SVD-LLM using their public codebases. While ASVD fails in certain cases due to numerical instability (denoted as NaN in the table), ERC-SVD consistently maintains stable and reliable performance. Additional zero-shot accuracy results comparison about these LLMs are provided in Table 9.
@@
ASVD 26.67 71.36 44.51
SVD-LLM 28.03 37.46 40.35
ERC-SVD 17.10 ($\downarrow$36%) 27.24 ($\downarrow$27%) 38.40 ($\downarrow$5%)
@@
ASVD 221.66 927.15 266.04
SVD-LLM 48.94 193.22 56.55
ERC-SVD 14.09 ($\downarrow$71%) 105.37 ($\downarrow$45%) 30.72 ($\downarrow$46%)
@@
ASVD 106.32 NaN NaN
SVD-LLM 12.42 104.27 39.55
ERC-SVD 11.57 ($\downarrow$7%) 69.28 ($\downarrow$34%) 27.24 ($\downarrow$31%)
Table 3: Perplexity ($\downarrow$) of larger-scale LLMs under 20% compression ratio on PTB [marcus1993building].
Method LLaMA-13B LLaMA-30B LLaMA-2-13B OPT-13B OPT-30B
SVD 1878.04 555.55 5464.57 1552.55 250.49
ASVD 12.42 31.36 81.00 36.47 30.95
SVD-LLM 12.17 9.10 88.13 14.86 12.94
ERC-SVD 9.70 8.41 66.47 13.22 12.89
Performance on larger-scale LLMs.
To examine the scalability of ERC-SVD, we evaluate its performance on larger-scale LLMs from two representative families: LLaMA and OPT (13B and 30B). As presented in Table 3, ERC-SVD consistently achieves superior performance over existing baselines under 20% compression ratio, demonstrating robust effectiveness across varying model scales. Additional zero-shot accuracy results about these larger-scale LLMs are provided in Table 6.
Table 4: Ablation results of REC and PLC on LLaMA-2-7B under 30% and 40% compression ratios.
[…]
30% ASVD - - NaN 0.15 0.27 0.51 0.26 0.22 0.53 0.20 0.31
SVD-LLM - - 34.96 0.21 0.42 0.55 0.34 0.22 0.60 0.23 0.37
ERC-SVD 30.68 0.22 0.42 0.58 0.35 0.23 0.61 0.24 0.38
24.77 0.22 0.48 0.60 0.38 0.25 0.65 0.22 0.40
23.28 0.23 0.51 0.62 0.42 0.29 0.68 0.24 0.43
Table 5: Perplexity ($\downarrow$) and zero-shot individual and average accuracy ($\uparrow$) of LLaMA-2-7B 30% compressed by ERC-SVD with different residual compensation factors ($\beta$).
[…]
0.025 10.38 0.23 0.51 0.62 0.41 0.28 0.67 0.23 0.42
0.050 10.32 0.23 0.51 0.62 0.42 0.29 0.68 0.24 0.43
0.075 10.11 0.22 0.49 0.61 0.39 0.28 0.66 0.23 0.41
0.010 10.46 0.23 0.47 0.60 0.38 0.26 0.66 0.23 0.40